What is the capacitance of a parallel plate capacitor with metal plates, each of area 1.00 m2, separated by 2.00 mm? What charge is stored in this capacitor if a voltage of 3.00 × 103 V is applied to it?


1
Expert's answer
2021-04-13T06:30:15-0400

The capacitance of a parallel plate capacitor is "C = \frac{\varepsilon \varepsilon_0 S}{d} = \frac{1 \cdot 8.85 \cdot 10^{-12} \frac{F}{m} \cdot 1 m^2 }{2 \cdot 10^{-3} m} \approx 4.4 nF".

The charge, stored in the capacitor is "q = C U = 4.4 nF \cdot 3 \cdot 10^3 V \approx 13.3 \cdot 10^{-6} C"